### v1

@@ Abstract
Multi-source information fusion (MSIF) leverages diverse data streams to enhance decision-making, situational awareness, and system resilience. Federated Learning (FL) enables MSIF while preserving privacy but suffers from client drift under high data heterogeneity, leading to performance degradation. Traditional mitigation strategies rely on reference-based gradient adjustments, which can be unstable in partial participation settings. To address this, we propose Gradient Centralized Federated Learning (GC-Fed), a reference-free gradient correction method inspired by Gradient Centralization (GC). We introduce Local GC and Global GC, applying GC during local training and global aggregation, respectively. Our hybrid GC-Fed approach selectively applies GC at the feature extraction layer locally and at the classifier layer globally, improving training stability and model performance. Theoretical analysis and empirical results demonstrate that GC-Fed mitigates client drift and achieves state-of-the-art accuracy gains of up to 20% in heterogeneous settings.
@@ 1 Introduction
Prior approaches to adjusting local training typically leverage historical information—such as the previous round’s global model, client updates, or control variates—to guide the next round optimization. These methods have proven effective in cross-silo settings (e.g., data center), where a majority of clients participate consistently. However, in cross-device scenarios (e.g., IoT devices and smartphones), where only a small fraction of thousands or millions of clients participate per round due to client availability and server hardware limitations, the sampled subset may fail to accurately reflect the optimal update direction [6, 21]. As illustrated in Figure 1(a), partial client participation can yield a global update that deviates from the one obtained via full participation. Moreover, our preliminary experiments (Figure 1(b)) demonstrate that the discrepancy between the true update (aggregated from all clients) and the partial update increases as participation diminishes. In other words, the historical information generated in previous rounds—even when not limited to the global model—can be less effective in guiding the current update when the participating clients do not reliably represent the entire population.
@@ 4.2.2 Small Scale Experiment
Table 5: Comparison of top-1 test accuracy (%) of FL algorithms across three datasets under a non-i.i.d. setting, with each dataset evaluated using its corresponding model.
Model (R) MLP (R: 200) CIFAR-10 CIFAR-10 CIFAR-100 CIFAR-100
Dataset EMNIST 5/100 5/200 5/100 5/200
Participation 5/1000
@@
As shown in Table 5, our proposed GC-Fed consistently outperforms standard baselines across various datasets and network architectures. For instance, on the EMNIST dataset using an MLP (with 5/1000 participation and LDA $\alpha=0.05$), GC-Fed achieves 66.63%, which is a +7.18 percentage point improvement over FedAvg (59.45%) and slightly surpasses FedVARP (66.18%). In the CNN experiments on CIFAR-10, the performance gains over FedAvg range from approximately +6.0 percentage points (e.g., 72.44% vs. 66.44% for 5/100 participation with $\alpha=0.10$) to as high as +19.1 percentage points (64.19% vs. 45.14% for 5/200 participation with $\alpha=0.05$). When compared against the best performing non–GC baseline (e.g., FedVARP or FedLC), the improvements vary between +3.5 and +7.3 percentage points. For CIFAR-100 with CNN, the improvements over FedAvg are in the range of approximately +6.9 to +8.2 percentage points, while the gains over the best non–GC baseline are more modest, spanning roughly +1.4 to +2.8 percentage points.

### v2

@@ Abstract
Federated Learning (FL) enables privacy-preserving multi-source information fusion (MSIF) but is challenged by client drift in highly heterogeneous data settings. Many existing drift-mitigation strategies rely on reference-based techniques—such as gradient adjustments or proximal loss—that use historical snapshots (e.g., past gradients or previous global models) as reference points. When only a subset of clients participates in each training round, these historical references may not accurately capture the overall data distribution, leading to unstable training. In contrast, our proposed Gradient Centralized Federated Learning (GC-Fed) employs a hyperplane as a historically independent reference point to guide local training and enhance inter-client alignment. GC-Fed comprises two complementary components: Local GC, which centralizes gradients during local training, and Global GC, which centralizes updates during server aggregation. In our hybrid design, Local GC is applied to feature-extraction layers to harmonize client contributions, while Global GC refines classifier layers to stabilize round-wise performance. Theoretical analysis and extensive experiments on benchmark FL tasks demonstrate that GC-Fed effectively mitigates client drift and achieves up to a 20% improvement in accuracy under heterogeneous and partial participation conditions.
@@ 1 Introduction
Prior approaches to adjusting local training typically leverage historical information—such as the previous round’s global model, client updates, or control variates—to guide the next round of optimization. These methods have proven effective in cross-silo settings (e.g., data centers), where a majority of clients participate consistently. However, in cross-device scenarios (e.g., IoT devices and smartphones), where only a small fraction of thousands or millions of clients participate per round due to client availability and server hardware limitations, the sampled subset may fail to accurately reflect the optimal update direction [6, 21]. As illustrated in Figure 1(a), partial client participation can yield a global update that deviates from the one obtained via full participation. Moreover, our preliminary experiments (Figure 1(b)) demonstrate that the discrepancy between the true update (aggregated from all clients) and the partial update increases as participation diminishes. In other words, the historical information generated in previous rounds—even when not limited to the global model—can be less effective in guiding the current update when the participating clients do not reliably represent the entire population.
@@ 4.2.2 Small Scale Experiment
Table 5: Comparison of top-1 test accuracy (%) of FL algorithms across three datasets under a non-i.i.d. setting. EMNIST is evaluated using an MLP for 200 rounds, while CIFAR-10 and CIFAR-100 are evaluated using a CNN for 800 rounds.
Dataset EMNIST CIFAR-10 CIFAR-100
Participation 5/1000 5/100 5/200 5/100 5/200
@@
As shown in Table 5, our proposed GC-Fed consistently outperforms standard baselines across various datasets and model architectures. For instance, on the EMNIST dataset using an MLP (with 5/1000 participation and LDA $\alpha=0.05$), GC-Fed achieves 66.63%, which is a +7.18 percentage point improvement over FedAvg (59.45%) and slightly surpasses FedVARP (66.18%). In the CNN experiments on CIFAR-10, the performance gains over FedAvg range from approximately +6.0 percentage points (e.g., 72.44% vs. 66.44% for 5/100 participation with $\alpha=0.10$) to as high as +19.1 percentage points (64.19% vs. 45.14% for 5/200 participation with $\alpha=0.05$). When compared against the best performing non–GC baseline (e.g., FedVARP or FedLC), the improvements vary between +3.5 and +7.3 percentage points. For CIFAR-100 with CNN, the improvements over FedAvg are in the range of approximately +6.9 to +8.2 percentage points, while the gains over the best non–GC baseline are more modest, spanning roughly +1.4 to +2.8 percentage points.